# An Eclipse-Ballooning Study of Shadow Bands During the April 2024 Total Eclipse

### v1

@@ 2.4 Observing Log
1The photodiode area for this study was 12.96 mm2, whereas it was 1.21 mm2 in our 2017 study[5], and our electronics were improved to minimize noise.
@@ 3.1 Spectrogram Analysis and Data Cleaning
Figures 13 illustrate this full cleaning process for HAB BL00. Raw and cleaned voltage series are shown with their corresponding spectrograms. These same results are essentially identical to those from three other HAB photodiodes (BL01–BL03), as shown in Appendix Figures A1A3. If shadow bands were present, they would appear as horizontal bands over a limited range of frequencies in the spectrograms. However, after cleaning, no persistent frequency signatures between 0.25 Hz and 100 Hz were detected from any of the HAB photodiodes, supporting the conclusion that shadow bands were absent at high altitudes over Concan.
Figure 1: Raw and cleaned light curve from the HAB BL00 photodiode sensor during the April 8, 2024, total solar eclipse. The original signal (blue) includes sharp voltage dips caused by the payload line intermittently blocking the photodiode sensor. The cleaned signal (black) preserves the eclipse structure while removing these interference artifacts. Time is shown in UTC.
Figure 2: Raw spectrogram of the HAB BL00 light intensity data during the April 8, 2024, total solar eclipse. Spectrogram computed using a Fast Fourier Transform (FFT) with a 5-second boxcar window and 50% overlap. The shown frequency range spans 0–25 Hz. Intensity is shown in decibels (dB). No shadow band signal was detected.
Figure 3: Cleaned spectrogram of HAB BL00 light intensity data during the April 8, 2024, total solar eclipse. The cleaning process partially removed voltage dips caused by intermittent payload line interference with the photodiode sensor. After cleaning, no shadow bands are detected. The spectrogram is free of distinct periodic signals up to 25 Hz (shown) and even up to 100 Hz (not shown).
@@ 3.2 Comparison of HAB, Aircraft, and Ground Data
After cleaning and generating the spectrograms, no periodic signals indicative of shadow bands were observed in the HAB BL00, BL01, BL02, and BL03 data (Figures 13, A1A3). These consistent results, which include four independent high altitude light sensors, confirm that our removal of interference artifacts did not introduce artificial features or mask real oscillations in the frequency domain. Moreover, as noted in Section 2.1, the sensitivity and noise characteristics of our photodiodes was improved relative to our similar 2017 study. The lack of any persistent signatures between 0.25 Hz and 100 Hz indicates the absence of shadow bands at stratospheric altitudes above Concan.
@@ 3.3 Atmospheric Turbulence and Shadow Bands
Thus, Figures 4 and 5, which show wind speed components and temperature profiles above Concan and New Hampshire, respectively, indicate that all optical measurements from the HABs and aircraft were collected well above the PBL, which isolated our searches for shadow bands from turbulence-induced variability within 3 km of the ground.
Figure 4: Wind speed components and temperature profiles from radiosonde flights over Concan, TX. Flight 01 (Top) and Flight 29 (Bottom) show that the PBL, characterized by temperature inversions and abrupt wind speed/direction changes, remains consistently below 3 km throughout all flights (all flights not shown). The vertical red dashed line highlights that the PBL is positioned between 1500 m and 2500 m. Flight 25b (Middle), which occurred during the total eclipse, exhibits a significant change in the north-south wind speed component in the same altitude range where a slight temperature inversion is observed, showing that the PBL is below 3 km during the eclipse.
Figure 5: Wind speed components and temperature profiles from radiosonde flights over New Hampshire. Flight 02 (Top) and Flight 28 (Bottom) show a consistently stable PBL just above 1000 m, as indicated by the vertical red dashed line marking the location of temperature inversions. Flight 25 (Middle), conducted closest to the time of the total eclipse, also shows a clear PBL inversion near this height as well. While no radiosonde was launched during totality, these flights provide a representative view of PBL conditions in the region. The stability of the PBL in these flights contrasts with the more variable PBL observed in Concan.

### v2

# arXiv:2601.13335v2[physics.ao-ph] 26 Feb 2026

@@ 2.4 Observing Log
1The photodiode area for this study was 12.96 mm2, 96 mm2, whereas a prior 2017 eclipse-ballooning study reported 1.21 mm2 [5], The electronics in the present study were improved to minimize noise.
@@ 3.1 Spectrogram Analysis and Data Cleaning
Figures $3.1$ illustrate this full cleaning process for HAB BL00. Raw and cleaned voltage series are shown with their corresponding spectrograms. These same results are essentially identical to those from three other HAB photodiodes (BL01–BL03), as shown in Appendix Figures A1A3. If shadow bands were present, they would appear as horizontal bands over a limited range of frequencies in the spectrograms. However, after cleaning, no persistent frequency signatures between 0.25 Hz and 100 Hz were detected from any of the HAB photodiodes, supporting the conclusion that shadow bands were absent at high altitudes over Concan.
Raw and cleaned light curve from the HAB BL00 photodiode sensor during the April 8, 2024, total solar eclipse. The original signal (blue) includes sharp voltage dips caused by the payload line intermittently blocking the photodiode sensor. The cleaned signal (black) preserves the eclipse structure while removing these interference artifacts. Time is shown in UTC.
Raw spectrogram of the HAB BL00 light intensity data during the April 8, 2024, total solar eclipse. Spectrogram computed using a Fast Fourier Transform (FFT) with a 5-second boxcar window and 50% overlap. The shown frequency range spans 0–25 Hz. Intensity is shown in decibels (dB). No shadow band signal was detected.
Cleaned spectrogram of HAB BL00 light intensity data during the April 8, 2024, total solar eclipse. The cleaning process partially removed voltage dips caused by intermittent payload line interference with the photodiode sensor. After cleaning, no shadow bands are detected. The spectrogram is free of distinct periodic signals up to 25 Hz (shown) and up to 100 Hz (not shown).
Figure 1: Illustration of the full cleaning and analysis process for the HAB BL00 photodiode sensor. (a) Raw and cleaned voltage time series showing removal of payload-line interference while preserving the eclipse morphology. (b) Corresponding raw spectrogram, dominated by nonphysical artifacts. (c) Cleaned spectrogram after artifact suppression, showing no persistent frequency signatures associated with shadow bands.
@@ 3.2 Comparison of HAB, Aircraft, and Ground Data
After cleaning and generating the spectrograms, no periodic signals indicative of shadow bands were observed in the HAB BL00, BL01, BL02, and BL03 data (Figures $3.1$, A1A3). These consistent results, which include four independent high altitude light sensors, confirm that our removal of interference artifacts did not introduce artificial features or mask real oscillations in the frequency domain. Moreover, as noted in Section 2.1, the sensitivity and noise characteristics of our photodiodes was improved relative to our similar 2017 study. The lack of any persistent signatures between 0.25 Hz and 100 Hz indicates the absence of shadow bands at stratospheric altitudes above Concan.
@@ 3.3 Atmospheric Turbulence and Shadow Bands
Thus, Figures 2 and 3, which show wind speed components and temperature profiles above Concan and New Hampshire, respectively, indicate that all optical measurements from the HABs and aircraft were collected well above the PBL, which isolated our searches for shadow bands from turbulence-induced variability within 3 km of the ground.
(a) Flight 01
(b) Flight 25b
(c) Flight 29
Figure 2: Wind speed components and temperature profiles from radiosonde flights over Concan, TX. Flight 01 (a) and Flight 29 (c) show that the PBL, characterized by temperature inversions and abrupt wind speed/direction changes, remains consistently below 3 km throughout all flights (all flights not shown). The vertical red dashed line highlights that the PBL is positioned between 1500 m and 2500 m. Flight 25b (b), which occurred during the total eclipse, exhibits a significant change in the north-south wind speed component in the same altitude range where a slight temperature inversion is observed, showing that the PBL is below 3 km during the eclipse.
(a) Flight 02
(b) Flight 25
(c) Flight 28
Figure 3: Wind speed components and temperature profiles from radiosonde flights over New Hampshire. Flight 02 (a) and Flight 28 (c) show a consistently stable PBL just above 1000 m, as indicated by the vertical red dashed line marking the location of temperature inversions. Flight 25 (b), conducted closest to the time of the total eclipse, also shows a clear PBL inversion near this height as well. While no radiosonde was launched during totality, these flights provide a representative view of PBL conditions in the region. The stability of the PBL in these flights contrasts with the more variable PBL observed in Concan.